OpenStudy (anonymous):

Algebra 2: Find the real or imaginary solutions of the equation by factoring: x^4 - 10x^2 + 16 = 0

OpenStudy (zarkon):

use the quadratic equation

OpenStudy (anonymous):

If you factor you will get \( \left(-8+x^2\right) \left(-2+x^2\right) \)

OpenStudy (anonymous):

would I have to use it with every polynomial that begins with a degree of 4?

OpenStudy (zarkon):

or factor :)

OpenStudy (zarkon):

\[u=x^2\]

OpenStudy (anonymous):

Right, I have issues with those problems that begin with x4.

OpenStudy (anonymous):

x^4*

OpenStudy (zarkon):

all 4 solutions are real

OpenStudy (anonymous):

oppps! mea culpa :(

OpenStudy (anonymous):

\( \left(x^2-8\right) \left(x^2-2\right) \implies x = \pm \sqrt 2 , \pm 2\sqrt{2} \)

OpenStudy (anonymous):

Sweet, thanks!

OpenStudy (anonymous):

Glad to help :)